# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v4

@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in (33). Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by 5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
Appendix G Proof of Theorem 5.4
@@
where we used 3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\mathscr{X}_{i}\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$, where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$, and $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in (17) and (21), respectively. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma G.1, Lemma G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
Note that we are retaining the dependence on $\mathscr{X}_{i}=\mathscr{X}_{i}(s_{i})$ to demonstrate the way that $s_{i}$ affects the rate of the regret. Substituting back the value $\tau\propto T^{\xi}$ into Appendix G, we get
@@
These bounds retain the constants $\mathscr{X}_{i}$ and $L_{\bm{\Gamma}}$ because they depend on the parameters $s_{1},\dots,s_{N}$ (or equivalently $c_{1},\dots,c_{N}$). Note that the dependence on $\{s_{j}\}_{j\in[N]}$ enters only through the multiplicative constant $\mathscr{X}_{i}$ and the additive constant $L_{\bm{\Gamma}}$, never through the powers of $T$. This makes explicit how the choice of $s_{i}$ influences the regret and the convergence to the NE. Below we show that if $s_{i}$ are the optimal-concave parameters (as defined in Definition L.3), i.e.
@@
    Multiplicative constants $\mathscr{X}_{i}$. The constant $\mathscr{X}_{i}=\{\mathscr{C}_{i},\mathscr{B}_{i}\}$ contains $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ (defined in (17)–(21)), which are the multiplicative constants in the concentration bounds for
@@
    respectively. Any value $s_{i}^{\prime}\neq s_{i}$ does not affect the rate of linear estimator $\widetilde{\bm{\theta}}_{i}$, but it could introduce a bias in the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. Precisely, if we choose $s_{i}^{\prime}\leq s_{i}$, the inclusion property of Proposition L.1 implies that an $s_{i}$-optimal-concave function remains $s_{i}^{\prime}$-concave. Thus no bias is introduced in the concentration of $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}$, whereas a choice $s_{i}^{\prime}>s_{i}$ would. Then for $s_{i}^{\prime}\leq s_{i}$, since $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are uniformly bounded in $s_{i}$, i.e. $\mathscr{B}_{i},\mathscr{C}_{i}<C$ for some $C>0$ independent of $s_{i}$, they do not affect the regret or NE rates.
@@
G.1 Proof of Lemma G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition D.3, and
@@
By Proposition D.3 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.
@@
Analyzing $A$ on $\mathcal{E}_{i}$: Define the event $\mathcal{Q}_{i}=\{\sup_{u\in\mathcal{U}}|\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(u)-\psi_{i,\widetilde{\bm{\theta}}_{i}}(u)|\leq\mathscr{C}_{i}\rho_{n_{i}^{(2)}}^{2/5}\}$. For $n_{i}^{(2)}$ sufficiently large, by Theorem D.6 we have that for every $\gamma>4$
@@
Analyzing $B$ on $\mathcal{E}_{i}$: By Proposition D.4, $\psi_{i,\widetilde{\bm{\theta}}_{i}}$ is Lipschitz, then
@@
Analyzing $C$ on $\mathcal{E}_{i}$: By Proposition D.4 we have
@@
which is strictly positive by 3.2. Then we have
@@ Appendix H NPLS for $s$-concave regression
where $\mathcal{C}$ is the set of all concave functions $\phi:\mathcal{U}\rightarrow[\underline{B}_{\phi},\bar{B}_{\phi}]$. Note that the problem is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$. When $h(x)=x$ (or equivalently for $1$-concave functions, i.e. concave functions), (41) was studied by dumbgen2004consistency, for which $\widehat{\phi}_{n}$ exists and is unique and doesn’t need restrictions on the codomain of $\phi_{0}$, that is $\underline{B}_{\phi},\bar{B}_{\phi}]$ can be the whole $\mathbb{R}$. Existence and uniqueness come from the coercivity of the loss $\mathcal{L}$, which is a property that only applies for $h(x)=x$. Since in our case $h$ is not necessarily the identity map, we rely on Weierstrass’ theorem (which guarantees the existence of at least a minimum for a continuous function $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$ defined on a compact set $\Omega$ of $\mathbb{R}^{n}$). Thus a solution of (41) exists (but might not be unique) because the $n$-dimensional variable $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$ where $\underline{B}_{\phi},\bar{B}_{\phi}]$ is compact. Uniqueness is not an issue in our case since we only care about the construction of a consistent estimator. Proposition H.1, which is followed by Lemma H.7, shows the constraints on the codomain of $\phi_{0}$ when $\psi_{0}:\mathcal{U}\to[\underline{B}_{\psi},\bar{B}_{\psi}]$ is $s$-concave for $s\neq 1$.
@@
Suppose that $\psi_{0}$ is $s$-concave, that is $\psi_{0}=h_{s}\circ\psi_{0}$ for some $s\neq 1$, where $h_{s}=d_{s}^{-1}$ as defined in (4). Then
@@
We now establish the assumptions to guarantee the convergence of the estimator defined in Equation 41. We consider a triangular scheme of observations $U_{i}=U_{n,i}$ and $Y_{i}=Y_{n,i}$ but suppress the additional subscript $n$ for notational simplicity. Let $\mathscr{M}_{n}$ be the empirical distribution of the design points $U_{i}$, i.e.
@@
H.2 holds if the density $f_{U}$ of the measure $\mathscr{M}$ is bounded away from zero, that is $f_{U}\geq C_{\mathscr{M}}$, which is a standard assumption for uniform convergence of non-parametric functions (see e.g. dumbgen2004consistency; mosching2020monotone) as well as for $L^{2}$ convergence (see e.g. groeneboom2018current; balabdaoui2019least).
@@
Suppose that H.2, H.3 and H.4 are satisfied. Let $\widehat{\phi}_{n}$ be a solution of (41), and let $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$. Then, for all $\gamma>4$ there exists a integer $n_{0}$ and a real value $\mathscr{C}>0$ such that
@@
The proof can be found in Section H.3.
@@
If $\psi_{0}$ and $\widehat{\psi}_{n}$ depends also on a parameter $\bm{\theta}\in\Theta$, $\psi_{0,\bm{\theta}}$ and $\widehat{\psi}_{n,\bm{\theta}}$, for which the domain $\mathcal{U}$ does not depend on $\bm{\theta}$ and such that H.2 and H.4 hold uniformly in $\bm{\theta}\in\Theta$ and all the constants $C_{\mathscr{M}},\bar{B}_{h},\underline{B}_{h},L,\alpha,\gamma,a,b$ are independent of $\bm{\theta}$, then
@@ H.2 Technical Lemma 2
H.2 implies the following Lemma H.8 that will be used to prove the uniform convergence consistency in Theorem H.5.
@@
Let $\rho_{n}\triangleq\frac{\log n}{n}$ and $\text{Leb}(\cdot)$ stands for Lebesgue measure and $U_{1},U_{2},\dots,U_{n}$ i.i.d. points with probability measure $M$ that satisfies H.2, then for a given constant $C_{3}>0$, and for any $\gamma>2$ there exists $n_{0}\in\mathbb{N}$ and a sequence $\epsilon_{n}=\epsilon_{n}(\gamma,C_{3},\alpha)>0$, $\epsilon_{n}\rightarrow 0$ such that
@@
H.3 Proof of Theorem H.5
First note that for every $\phi\in\mathcal{C}$, since $\phi\in[\underline{B}_{\phi},\bar{B}_{\phi}]$, then $h\circ\phi\in[\underline{B}_{\psi},\bar{B}_{\psi}]$. By assumption we also have $h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ for some $0<\underline{B}_{h}<\bar{B}_{h}<\infty$. Let $\widehat{\phi}_{n}$ be a solution of (41), i.e. the LS estimate of $\phi_{0}$. Then, being $\widehat{\phi}_{n}$ a minimizer, we have that for any direction $\Delta$ such that $\widehat{\phi}_{n}+\delta\Delta\in\mathcal{C}$ for $\delta\geq 0$ sufficiently small, $\mathcal{L}(\widehat{\phi}_{n}+\delta\Delta)\geq\mathcal{L}(\widehat{\phi}_{n})$. Hence the directional derivative at $\widehat{\phi}_{n}$ along direction $\Delta$ satisfies
@@
In what follows, we apply (44) to a special class of perturbation functions $\Delta$ and write
@@
The next Lemma H.10 is proved in Section H.4
@@
The next ingredient for our proof is a result concerning the difference of two concave functions, which can be found in dumbgen2004consistency and we report the proof in Section H.5 for completeness.
@@
Let $\phi$ satisfy H.4. There is a universal constant $\widetilde{\mathscr{C}}=\widetilde{\mathscr{C}}(L,\alpha)$ with the following property: for any $\xi>0$, let $\delta\leq\widetilde{\mathscr{C}}\min\left(b-a,\xi^{1/\alpha}\right)$. Then, for any concave function $\hat{\phi}$
@@
Now, we have to show that one of our classes $\mathcal{D}_{m}$ does indeed contain useful perturbation functions $\Delta$. We denote with $\check{\phi}$ the unique continuous and piecewise linear function with knots in $\mathscr{T}\cap\left(U_{1},U_{n}\right)$ such that $\check{\phi}_{n}=\widehat{\phi}_{n}$ on $\mathscr{T}=\{U_{1},U_{2},\dots,U_{n}\}$. Thus $\check{\phi}_{n}$ is one particular LS estimator for $\phi_{0}$. The following technical Lemma H.12 can be found in dumbgen2004consistency.
@@
From Lemma H.11, replacing
@@
Because $c_{n},c_{n}+\delta_{n}]$ is a sub-interval of $\mathcal{U}$ on length $\delta_{n}$, Lemma H.8 (which holds by H.2) guarantees that this interval contains at least one observation $U_{i}$. This implies Lemma H.12. For any $\Delta\in\mathcal{D}_{6}$ define $I_{\Delta}=\{i:\Delta(U_{i})<0\}$. Define also $J=\{i:\phi_{0}(U_{i})-\widehat{\phi}_{n}(U_{i})>0\}$. Using that $0<h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ with $\underline{B}_{h},\bar{B}_{h}>0$ (which comes from Lemma H.7) we have that
@@
Recall that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$. By the definition of $S_{n}(6)$ and Lemma H.12, there is a (random) function $\Delta\in\mathcal{D}_{6}$ such that
@@
Now consider the event $B_{6,n,m}$ in Lemma H.10 and the event $A_{n,\gamma}$ in Lemma H.8. For $\gamma>4$ define
@@
then with probability less than $2n^{2-\gamma}$ we have $\sup_{\left[a+\delta_{n},b-\delta_{n}\right]}|\psi_{0}-\widehat{\psi}_{n}|\geq C_{\mathscr{C}}\gamma\delta_{n}^{\alpha}$ provided (by Lemma H.10)
@@
More precisely, the condition on $\mathscr{C}$ is that $\mathscr{C}\geq\max\left\{\widetilde{\mathscr{C}}^{-\alpha}\bar{B}_{h},\frac{\gamma\sigma\sqrt{7}}{(4\bar{B}_{h})^{-1}\|o_{h}\|^{-1}\underline{B}_{h}^{2}\left(C_{\mathscr{M}}/8\right)^{1/2}}\right\}$, where $\gamma>4$. Using that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$ and $\widetilde{\mathscr{C}}^{\alpha}=\frac{\alpha}{4^{1+\alpha}L}$ (from Lemma H.11), this condition becomes
@@
H.4 Proof of Lemma H.10
H.3 implies the following inequality
@@
This defines a collection $\Phi$ of at most $n^{2}$ different nonzero functions $\zeta_{jk}^{(e)}$. Then for any fixed $\gamma_{o}>2$, (47) implies

### v5

@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in equation 33. Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by ˜5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
Appendix G Proof of Theorem˜5.4
@@
where we used ˜3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\mathscr{X}_{i}\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$, where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$, and $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in equation 17 and equation 21, respectively. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma˜G.1, Lemma˜G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
Note that we are retaining the dependence on $\mathscr{X}_{i}=\mathscr{X}_{i}(s_{i})$ to demonstrate the way that $s_{i}$ affects the rate of the regret. Substituting back the value $\tau\propto T^{\xi}$ into Appendix˜G, we get
@@
These bounds retain the constants $\mathscr{X}_{i}$ and $L_{\bm{\Gamma}}$ because they depend on the parameters $s_{1},\dots,s_{N}$ (or equivalently $c_{1},\dots,c_{N}$). Note that the dependence on $\{s_{j}\}_{j\in[N]}$ enters only through the multiplicative constant $\mathscr{X}_{i}$ and the additive constant $L_{\bm{\Gamma}}$, never through the powers of $T$. This makes explicit how the choice of $s_{i}$ influences the regret and the convergence to the NE. Below we show that if $s_{i}$ are the optimal-concave parameters (as defined in Definition˜L.3), i.e.
@@
    Multiplicative constants $\mathscr{X}_{i}$. The constant $\mathscr{X}_{i}=\{\mathscr{C}_{i},\mathscr{B}_{i}\}$ contains $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ (defined in equation 17–equation 21), which are the multiplicative constants in the concentration bounds for
@@
    respectively. Any value $s_{i}^{\prime}\neq s_{i}$ does not affect the rate of linear estimator $\widetilde{\bm{\theta}}_{i}$, but it could introduce a bias in the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. Precisely, if we choose $s_{i}^{\prime}\leq s_{i}$, the inclusion property of Proposition˜L.1 implies that an $s_{i}$-optimal-concave function remains $s_{i}^{\prime}$-concave. Thus no bias is introduced in the concentration of $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}$, whereas a choice $s_{i}^{\prime}>s_{i}$ would. Then for $s_{i}^{\prime}\leq s_{i}$, since $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are uniformly bounded in $s_{i}$, i.e. $\mathscr{B}_{i},\mathscr{C}_{i}<C$ for some $C>0$ independent of $s_{i}$, they do not affect the regret or NE rates.
@@
G.1 Proof of Lemma˜G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition˜D.3, and
@@
By Proposition˜D.3 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.
@@
Analyzing $A$ on $\mathcal{E}_{i}$: Define the event $\mathcal{Q}_{i}=\{\sup_{u\in\mathcal{U}}|\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(u)-\psi_{i,\widetilde{\bm{\theta}}_{i}}(u)|\leq\mathscr{C}_{i}\rho_{n_{i}^{(2)}}^{2/5}\}$. For $n_{i}^{(2)}$ sufficiently large, by Theorem˜D.6 we have that for every $\gamma>4$
@@
Analyzing $B$ on $\mathcal{E}_{i}$: By Proposition˜D.4, $\psi_{i,\widetilde{\bm{\theta}}_{i}}$ is Lipschitz, then
@@
Analyzing $C$ on $\mathcal{E}_{i}$: By Proposition˜D.4 we have
@@
which is strictly positive by ˜3.2. Then we have
@@ Appendix H NPLS for $s$-concave regression
where $\mathcal{C}$ is the set of all concave functions $\phi:\mathcal{U}\rightarrow[\underline{B}_{\phi},\bar{B}_{\phi}]$. Note that the problem is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$. When $h(x)=x$ (or equivalently for $1$-concave functions, i.e. concave functions), equation 41 was studied by dumbgen2004consistency, for which $\widehat{\phi}_{n}$ exists and is unique and doesn’t need restrictions on the codomain of $\phi_{0}$, that is $\underline{B}_{\phi},\bar{B}_{\phi}]$ can be the whole $\mathbb{R}$. Existence and uniqueness come from the coercivity of the loss $\mathcal{L}$, which is a property that only applies for $h(x)=x$. Since in our case $h$ is not necessarily the identity map, we rely on Weierstrass’ theorem (which guarantees the existence of at least a minimum for a continuous function $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$ defined on a compact set $\Omega$ of $\mathbb{R}^{n}$). Thus a solution of equation 41 exists (but might not be unique) because the $n$-dimensional variable $(\phi(U_{1}),\phi(U_{2}),\dots,\phi(U_{n}))\in[\underline{B}_{\phi},\bar{B}_{\phi}]$ where $\underline{B}_{\phi},\bar{B}_{\phi}]$ is compact. Uniqueness is not an issue in our case since we only care about the construction of a consistent estimator. Proposition˜H.1, which is followed by Lemma˜H.7, shows the constraints on the codomain of $\phi_{0}$ when $\psi_{0}:\mathcal{U}\to[\underline{B}_{\psi},\bar{B}_{\psi}]$ is $s$-concave for $s\neq 1$.
@@
Suppose that $\psi_{0}$ is $s$-concave, that is $\psi_{0}=h_{s}\circ\psi_{0}$ for some $s\neq 1$, where $h_{s}=d_{s}^{-1}$ as defined in equation 4. Then
@@
We now establish the assumptions to guarantee the convergence of the estimator defined in Equation˜41. We consider a triangular scheme of observations $U_{i}=U_{n,i}$ and $Y_{i}=Y_{n,i}$ but suppress the additional subscript $n$ for notational simplicity. Let $\mathscr{M}_{n}$ be the empirical distribution of the design points $U_{i}$, i.e.
@@
˜H.2 holds if the density $f_{U}$ of the measure $\mathscr{M}$ is bounded away from zero, that is $f_{U}\geq C_{\mathscr{M}}$, which is a standard assumption for uniform convergence of non-parametric functions (see e.g. dumbgen2004consistency; mosching2020monotone) as well as for $L^{2}$ convergence (see e.g. groeneboom2018current; balabdaoui2019least).
@@
Suppose that ˜H.2, ˜H.3 and ˜H.4 are satisfied. Let $\widehat{\phi}_{n}$ be a solution of equation 41, and let $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$. Then, for all $\gamma>4$ there exists a integer $n_{0}$ and a real value $\mathscr{C}>0$ such that
@@
The proof can be found in Section˜H.3.
@@
If $\psi_{0}$ and $\widehat{\psi}_{n}$ depends also on a parameter $\bm{\theta}\in\Theta$, $\psi_{0,\bm{\theta}}$ and $\widehat{\psi}_{n,\bm{\theta}}$, for which the domain $\mathcal{U}$ does not depend on $\bm{\theta}$ and such that ˜H.2 and ˜H.4 hold uniformly in $\bm{\theta}\in\Theta$ and all the constants $C_{\mathscr{M}},\bar{B}_{h},\underline{B}_{h},L,\alpha,\gamma,a,b$ are independent of $\bm{\theta}$, then
@@ H.2 Technical Lemma 2
˜H.2 implies the following Lemma˜H.8 that will be used to prove the uniform convergence consistency in Theorem˜H.5.
@@
Let $\rho_{n}\triangleq\frac{\log n}{n}$ and $\text{Leb}(\cdot)$ stands for Lebesgue measure and $U_{1},U_{2},\dots,U_{n}$ i.i.d. points with probability measure $M$ that satisfies ˜H.2, then for a given constant $C_{3}>0$, and for any $\gamma>2$ there exists $n_{0}\in\mathbb{N}$ and a sequence $\epsilon_{n}=\epsilon_{n}(\gamma,C_{3},\alpha)>0$, $\epsilon_{n}\rightarrow 0$ such that
@@
H.3 Proof of Theorem˜H.5
First note that for every $\phi\in\mathcal{C}$, since $\phi\in[\underline{B}_{\phi},\bar{B}_{\phi}]$, then $h\circ\phi\in[\underline{B}_{\psi},\bar{B}_{\psi}]$. By assumption we also have $h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ for some $0<\underline{B}_{h}<\bar{B}_{h}<\infty$. Let $\widehat{\phi}_{n}$ be a solution of equation 41, i.e. the LS estimate of $\phi_{0}$. Then, being $\widehat{\phi}_{n}$ a minimizer, we have that for any direction $\Delta$ such that $\widehat{\phi}_{n}+\delta\Delta\in\mathcal{C}$ for $\delta\geq 0$ sufficiently small, $\mathcal{L}(\widehat{\phi}_{n}+\delta\Delta)\geq\mathcal{L}(\widehat{\phi}_{n})$. Hence the directional derivative at $\widehat{\phi}_{n}$ along direction $\Delta$ satisfies
@@
In what follows, we apply equation 44 to a special class of perturbation functions $\Delta$ and write
@@
The next Lemma˜H.10 is proved in Section˜H.4
@@
The next ingredient for our proof is a result concerning the difference of two concave functions, which can be found in dumbgen2004consistency and we report the proof in Section˜H.5 for completeness.
@@
Let $\phi$ satisfy ˜H.4. There is a universal constant $\widetilde{\mathscr{C}}=\widetilde{\mathscr{C}}(L,\alpha)$ with the following property: for any $\xi>0$, let $\delta\leq\widetilde{\mathscr{C}}\min\left(b-a,\xi^{1/\alpha}\right)$. Then, for any concave function $\hat{\phi}$
@@
Now, we have to show that one of our classes $\mathcal{D}_{m}$ does indeed contain useful perturbation functions $\Delta$. We denote with $\check{\phi}$ the unique continuous and piecewise linear function with knots in $\mathscr{T}\cap\left(U_{1},U_{n}\right)$ such that $\check{\phi}_{n}=\widehat{\phi}_{n}$ on $\mathscr{T}=\{U_{1},U_{2},\dots,U_{n}\}$. Thus $\check{\phi}_{n}$ is one particular LS estimator for $\phi_{0}$. The following technical Lemma˜H.12 can be found in dumbgen2004consistency.
@@
From Lemma˜H.11, replacing
@@
Because $c_{n},c_{n}+\delta_{n}]$ is a sub-interval of $\mathcal{U}$ on length $\delta_{n}$, Lemma˜H.8 (which holds by ˜H.2) guarantees that this interval contains at least one observation $U_{i}$. This implies Lemma˜H.12. For any $\Delta\in\mathcal{D}_{6}$ define $I_{\Delta}=\{i:\Delta(U_{i})<0\}$. Define also $J=\{i:\phi_{0}(U_{i})-\widehat{\phi}_{n}(U_{i})>0\}$. Using that $0<h^{\prime}\in[\underline{B}_{h},\bar{B}_{h}]$ with $\underline{B}_{h},\bar{B}_{h}>0$ (which comes from Lemma˜H.7) we have that
@@
Recall that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$. By the definition of $S_{n}(6)$ and Lemma˜H.12, there is a (random) function $\Delta\in\mathcal{D}_{6}$ such that
@@
Now consider the event $B_{6,n,m}$ in Lemma˜H.10 and the event $A_{n,\gamma}$ in Lemma˜H.8. For $\gamma>4$ define
@@
then with probability less than $2n^{2-\gamma}$ we have $\sup_{\left[a+\delta_{n},b-\delta_{n}\right]}|\psi_{0}-\widehat{\psi}_{n}|\geq C_{\mathscr{C}}\gamma\delta_{n}^{\alpha}$ provided (by Lemma˜H.10)
@@
More precisely, the condition on $\mathscr{C}$ is that $\mathscr{C}\geq\max\left\{\widetilde{\mathscr{C}}^{-\alpha}\bar{B}_{h},\frac{\gamma\sigma\sqrt{7}}{(4\bar{B}_{h})^{-1}\|o_{h}\|^{-1}\underline{B}_{h}^{2}\left(C_{\mathscr{M}}/8\right)^{1/2}}\right\}$, where $\gamma>4$. Using that $o_{h}=(\underline{B}_{h},\bar{B}_{h})^{\top}$ and $\widetilde{\mathscr{C}}^{\alpha}=\frac{\alpha}{4^{1+\alpha}L}$ (from Lemma˜H.11), this condition becomes
@@
H.4 Proof of Lemma˜H.10
˜H.3 implies the following inequality
@@
This defines a collection $\Phi$ of at most $n^{2}$ different nonzero functions $\zeta_{jk}^{(e)}$. Then for any fixed $\gamma_{o}>2$, equation 47 implies